find the two repeated limits of the function f(x,y)=(y-x/y+x)(1+x^2/1+y^2) at (0,0).Does the simultaneous limit of f exist as (x,y) tends to (0,0)?
1
Expert's answer
2020-04-07T12:38:09-0400

The given function is


"f(x,y)=\frac{(y-x)}{(y+x)} ×\frac{(1+x^2)}{(1+y^2)}"


The repeated limit of "f(x,y)" are following

"\lim_{x\to 0} \lim_{y\to 0} f(x,y)=\lim_{x\to 0}(\lim_{y\to 0} f(x,y))"

"=\lim_{x\to 0}(\lim_{y\to 0} \frac{(y-x)}{(y+x)}×\frac{(1+x^2)}{(1+y^2)})"

"=\lim_{x\to 0} (-1) (1+x^2)"

"=-1."

Again, "lim_{y\to 0}lim_{x\to 0}f(x,y)=\lim_{y\to 0}(\lim_{x\to 0} f(x,y))"

"=\lim_{y\to0}(\lim_{x\to0} \frac{(y-x)}{(y+x)}×\frac{(1+x^2)}{(1+y^2)} )"

"=\lim_{y\to0}\frac{1}{1+y^2}"

"=1"

Hence ,the repeated limits exist and unequal.

Now we have to find the simultaneous limit of "f(x,y)"


"\lim_{(x,y)\to (0,0)}f(x,y)=\lim_{(x,y)\to(0,0)}\frac{(y-x)}{(y+x)}×\frac{(1+x^2)}{(1+y^2)}"


Let "(x,y)\rightarrow(0,0)" along the line "y=mx" ,"m\in\R" then


"\lim_{(x,y)\to(0,0)}f(x,y)=\lim_{x\to0} \frac{(mx-x)}{(mx+x)} ×\frac{(1+x^2)}{(1+m^2x^2)}"




"=\frac{(1-m)}{(1+m)}×\lim_{x\to0} \frac{(1+x^2)}{(1+m^2x^2)}""=\frac{(1-m)}{(1+m)}"

Which is not unique as it takes different values for different values of "m" .

Hence the simultaneous limit does not exist.